The speed of sound v in a medium depend on it's wave length,the Young modulus ,density of the medium use the method of dimensional analysis to derive a formula for speed of sound in a medium.
E=kgm-1s-2
1
Expert's answer
2020-10-01T10:07:53-0400

The Dimension o f Young Modulus E is shown as;

"\frac{N}{m^2}=\frac{kgm}{s^2m^2}=\frac{kg}{s^2m}"


This means that;

"(\frac{kg}{s^2m})^\alpha (\frac{kg}{m^3})\beta (m)^\gamma=\frac{m}{s}"


Hence; "\alpha=0.5=-\beta and \gamma=0"

Consequently;


"v=\sqrt\frac{E}{\rho}"

i.e "\rho" is the Density


Answer: "v=\sqrt\frac{E}{\rho}"